OpenStudy (anonymous):

The legs of a right triangle measure 5 inches and 7 units. If (theta) is the angle between the 7-inch leg and the hypotenuse, sin(theta) = _____.

OpenStudy (campbell_st):

start by drawing a picture 

OpenStudy (anonymous):

\[\sin(\theta)= 7/hyp\] \[5^{2}+7^{2}=c ^{2}\] \[25+49=c ^{2}\] \[74=c ^{2}\] \[0.86023=c\] right?

OpenStudy (anonymous):

So \[\sin(\theta)=7/0.86023\] \[\sin(\theta)=8.14\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you have an error with the hypotenuse \[c^2 = 74...... c = \sqrt{74} .......... c = 8.6\] then find the side that is opposite the angle.... (its not 7) then write the ratio.....

OpenStudy (campbell_st):

remember, the hypotenuse is always the longest side....

OpenStudy (anonymous):

Oops, I had that written on my paper just typed it wrong.

OpenStudy (anonymous):

So, \[\sin(\theta)=7/8.6\] \[\sin(\theta)=0.81\]